Plz plz help me.i’m so confused……it’s urgent
 
NCERT question ;
 
On the basis of quantum numbers, justify that the sixth period of the periodic table
should have 32 elements.??
 
Plzzzz help me…I’m so confused..it’s urgent…plz
 Here principal quantum no (n)= 6
So the azimuthal quantum no (l)= (0 to n-1 i.e 0 to 5) 0,1,2,3,4,5,6
But  the value of l decides the subshells…
So here, for l=0 , subshell- s , no. of orbitals (m) =1 (i.e 0)
              for l=1 , subshell- p , no. of orbitals (m) =3 (i.e  -1,0,1)
                      for l=2 , subshell- d , no. of orbitals (m) =5 (i.e -2,-1,0,1,2)
               for l=3 , subshell- f , no. of orbitals (m) =7(i.e -3,-2,-1,0,1,2,3)
for l=4 , subshell- g , no. of orbitals (m) =9 (i.e -4,-3,-2,-1,0,1,2,3,4)
for l=5 , subshell- h , no. of orbitals (m) =11(i.e -5,-4,-3,-2,-1,0,1,2,3,4,5)
 
so the 6th shell includes g & h subshells also , but in 6th period we don’t have g or h subshells…so how can we ignore the total azimuthal quantum no. assigned for the principal quantum no. 6 ??
 
 
 
 
 

Grade:12th pass

1 Answers

Aarti Gupta
9 years ago
For ,n(principal quantum no.) = 6
The values of azimuthal quantum no. will be,l = 0,1,2,3,4,5
and not 6 as value of l = 0 to n-1
now sixth period starts with the filling of 6s orbital (n = 6).There will be sixteen orbitals which are-
One 6s,seven 4f, five 5d,and three 6p having
6s = 2e-
4f = 14 e-
5d = 10 e-
6p = 6e-
Thus total will be of 32 electrons,as such there are 32 elements in sixth period.